I've been trying to simplify the following expression and keep missing a step while doing so. Can you please assist me? Thank you!

1/2(3x + 2)- 1/4(3x - 2)=
[{{(1/2)*3x} + {(1/2)*(2/1)}} + {-1{{(1/4)*3x}-{(1/4)*2}}}]=
[{(3x/2)+(2/2)} + {-1{(3x/4)-(2/4)}}]=
[{(3x/2+1)} + {-1*(3x/4)}-{(-1)*(1/2)}}]=
[(3x/2+1)} + {{(-1)*(3x/4)}-{(-1)*(1/2)}]=
[{3x/2 + 1} + {-3x/4 + 1/2}]=
[{3x/2 + 1} + {-3x/4 +1/2}]=
[3x/2 +1 -3x/4 +1/2]=
3x/2 -3x/4 + 1+ (1/2)=
2/2*3x/2 - 3x/4 +1 1/2 =
(6x/4-3x/4) + 1.5 =
3x/4 +1.5
or 3/4x + 1.5

Found 2 solutions by stanbon, chiexpert:
Answer by stanbon(75887)

(1/2)(3x + 2)- (1/4)(3x - 2)
---
= (3/2)x + 1 -(3/4)x + (1/2)
----------------------------------
= (6/4)x + 1 -(3/4)x + (1/2)
---
= (3/4)x + (3/2)
======================
Cheers,
stan H.

Answer by chiexpert(48)

1/2(3x + 2)- 1/4(3x - 2)=
[{{(1/2)*3x} + {(1/2)*(2/1)}} + {-1{{(1/4)*3x}-{(1/4)*2}}}]= expand the equation
remember that subtracting a whole equation is same thing as adding a negative one in each when expanding. keep the subtract sign or negative consistant thruout the equations
[{(3x/2)+(2/2)} + {-1{(3x/4)-(2/4)}}]= same things here. now multiply out the fractions. keeps all sign consistent

[{(3x/2+1)} + {-1*(3x/4)}-{(-1)*(1/2)}}]=now the negative one is multiply out


[(3x/2+1)} + {{(-1)*(3x/4)}-{(-1)*(1/2)}]=
[{3x/2 + 1} + {-3x/4 + 1/2}]=
[{3x/2 + 1} + {-3x/4 +1/2}]=simplify
[3x/2 +1 -3x/4 +1/2]=now consolidate like equations

3x/2 -3x/4 + 1+ (1/2)=
2/2*3x/2 - 3x/4 +1 + 1/2 =simply further
(6x/4-3x/4) + 1.5 =
3x/4 +1 1/2 or 3/4x + 1.5